The Wongs bought a new house three years ago for $92000. The house is now worth about $113,000. Assuming a steady annual percentage growth rate, approximately what was the yearly rate of appreciation?
Answer by richwmiller(17219)

113,000-92000=21000
21000/3=about 7000 per annum
i=prt
21000=92000*r*3
r=0.076087=7.6087%